# Efficient Long Sequential Low-rank Adaptive Attention for Click-through rate Prediction

### v1

@@ 1. Introduction
In addressing the issue of long sequences, the industry commonly adopts information filtering techniques to select top-k items, often based on item attributes like SIM(Pi et al., 2020), or some form of low-precision attention like ETA(Chen et al., 2021). and SDIN(Cao et al., 2022) This approach has been validated in various studies (Chang et al., 2023; Si et al., 2024; Chen et al., 2021; Cao et al., 2022; Pi et al., 2020), including UBR4CTR (Qin et al., 2020), which employs end-to-end retrieval-augmented transformers, as well as TWIN (Chang et al., 2023) and TWIN V2 (Si et al., 2024), which utilize GPU-optimized hierarchical attention to process sequences and significantly enhance CTR prediction accuracy effectively. However, this approach sacrifices some products with low correlation in some aspects. It also causes information loss. For example, when the target item is pants, the system retrieves the user’s past behaviors related to buying pants generating the top K relevant features. However, it may overlook the user’s purchases of other light-colored clothes, which implies an interest in dark-colored pants.

### v2

@@ 1. Introduction
In addressing the issue of long sequences, the industry commonly adopts information filtering techniques to select top-k items, often based on item attributes like SIM(Pi et al., 2020), or some form of low-precision attention like ETA(Chen et al., 2021). and SDIM(Cao et al., 2022) This approach has been validated in various studies (Chang et al., 2023; Si et al., 2024; Chen et al., 2021; Cao et al., 2022; Pi et al., 2020), including UBR4CTR (Qin et al., 2020), which employs end-to-end retrieval-augmented transformers, as well as TWIN (Chang et al., 2023) and TWIN V2 (Si et al., 2024), which utilize GPU-optimized hierarchical attention to process sequences and significantly enhance CTR prediction accuracy effectively. However, this approach sacrifices some products with low correlation in some aspects. It also causes information loss. For example, when the target item is pants, the system retrieves the user’s past behaviors related to buying pants generating the top K relevant features. However, it may overlook the user’s purchases of other light-colored clothes, which implies an interest in dark-colored pants.